OpenStudy (anonymous):

suppose that x1=-1; x2=2; x3=4; x4=-3 is a solution of a nonhomogeneous linear system Ax=b and that the solution set of the homogeneous system Ax=0 given by the formulas x1=-3r+4s; x2=r-s; x3=r; x4=s. a) Find a vector form of the general solution of Ax =o b) Find a vector form of the general solution of Ax =b Anyone helps me, please

OpenStudy (sirm3d):

the vector form of the general solution to the homogeneous system AX = 0 is \[\left[\begin{matrix}-3r+4s\\r-s\\r\\s\end{matrix}\right]=r\left[\begin{matrix}-3\\1\\1\\0\end{matrix}\right]+s\left[\begin{matrix}4\\-1\\0\\1\end{matrix}\right]\]

OpenStudy (anonymous):

yes. thanks a lot. I got it. don't we have to form A? from those solution of x to get b and then get b)

OpenStudy (sirm3d):

i am thinking of this, but i don't know yet how to proceed. i'm not even sure if this is the right approach.\[AX=B \Rightarrow AX = 0 + B\]